OpenStudy (anonymous):

A researcher tagged 40 giant turtles in 2008. In 2009, he returned and captured 50 turtles, 16 of which were tagged. In 2010, she returned again and captured 30 turtles, 8 of which were tagged. A.) Use a proportion to estimate the turtle population of 2009 B.) Use a proportion to estimate the turtle population for 2010

OpenStudy (kirbykirby):

\[ Total = \frac{\text{original number tagged} \times \text{total recaptured}}{\text{number tagged on recapture}}\]

OpenStudy (kropot72):

A) The tagged turtles are 16/50 of the sample. Therefore an estimate of the 2009 population is: \[40\times\frac{50}{16}=you\ can\ calculate\]

OpenStudy (anonymous):

125.?

OpenStudy (kropot72):

Correct. Can you calculate B) with the same method?

OpenStudy (anonymous):

sweet thank you :)

OpenStudy (kropot72):

You're welcome :)